OpenStudy (anonymous):

If f(x)= 3x+5 And g(x)=x^2+1 what is the value of g{f(x)} A) 9x^2+30+26 B) 3x^3+5x^2+3x+5 C) 9x^2+25 D) 3x^2+8 I know it's between b and d Is it asking for g composition f?

OpenStudy (dayakar):

f(x)=3x+5---------(1) g(x)= x^2+1--------(2) g[f(x)] = g[ 3x+5] { from (1)}

OpenStudy (dayakar):

replace x with 3x+5 in equation (2) can u do it

OpenStudy (anonymous):

So do I multiply the two together and foil them or do I used the combination process and replace g(x) with f?

OpenStudy (dayakar):

just u take equation (2) both sides u plug in 3x+5 in the place of x do u understand

OpenStudy (anonymous):

So it will be (3x+5)^2+1

OpenStudy (dayakar):

g(3x+5) = (3x+5)^2 +1 simplify it u can get the answer

OpenStudy (anonymous):

Oh ok I understand now

OpenStudy (dayakar):

can u show ur work

OpenStudy (anonymous):

I got 9x^2+26

OpenStudy (dayakar):

(3x)^2+2*3x*5+5^2+1 = 9x^2+30x+25+1 =9x^2+30x+26